OpenStudy (anonymous):

how do you find the roots of y=x(squared)+2x=3 by factoring?

OpenStudy (anonymous):

do you know two numbers a and b, where a+b=2 and a*b=-3

OpenStudy (anonymous):

numbers 3 and -1 do result in that way.

OpenStudy (anonymous):

so you can write :x2+2x-3=(x-3)(x+1)=0

OpenStudy (anonymous):

so the answers are 3 and -1